OpenStudy (anonymous):

save this system of equation using substitution put your answer in ordered pair form, (x,Y) 2x+6y=32, 5x-2y=12

OpenStudy (anonymous):

alright so from your two equations i will label that as (1) 2x+6y=32, and (2) 5x-2y=12 from (1) let us solve for either x or y i will just solve for s and then sub it in (2) so therefore from (1) i got 2x = 32 - 6y and --> x=16-3y (3) and now i will solve this equation in (2) to obtain 5(16-3y) - 2y =12 and now i will just solve for y to obtain 80 - 15y -2y =12 -->68 = 17y and y=68/17 = 4 therefore y=4 and now letus finally sub it in (3) to get x= 16 - 3(4) --> 16 - 12 = 4 therefore the answer is (4,4)

OpenStudy (anonymous):

how do you set in in X,Y Format?

OpenStudy (anonymous):

if x=4 then y must =4 as well by our equations